# BACK TO THE FUTURE CYCLOPEAN STEREO: A Human Perception Approach Unifying Deep and Geometric Constraints

### v1

@@ 1 Introduction
What does it mean to “understand stereo vision” and why does it matter? Understanding means discovering internal representations of the problem at hand, and often it is needed to generalize to new situations or to allow us to apply such knowledge broadly. Suppose that a camera records video data of objects falling to the ground. It is plausible that an algorithm for machine learning (ML) could make predictions of such trajectories with an accuracy equal to or greater than any physics model, assuming enough variety and quantity of data are provided. However, such an ML solution would not have revealed the underlying laws of motion that govern them, such as the formula $F=ma$ and $a=9.8m/s^{2}$, an abstraction of reality that alone does not even give the most accurate models when the wind plays a role in the trajectory. Still, it has been a key insight for engineering and our civilization. Our goal is to better understand stereo vision with the mandate that to demonstrate better understanding the new solution must perform on par or better with respect to the state-of-the-art stereo vision. In this way, the abstractions of the 3D reality we make here can help not only to improve efficiency and performance of stereo algorithms, but perhaps such 3D models of reality can be used for other activities in computer vision.
@@ 1.2 Our Contribution
Cyclopean Eye Model: a yet novel framework that handles depth discontinuities and occlusions, where for opaque surfaces there is one and only one disparity solution per cyclopean spatial coordinate $(e,x)$, allowing for a more human-like perception of 3D scenes (see Figures 2, 3, 4).
Occlusion and Texture-Less Region Recovery: incorporate a prior monocular surface model to fill in occluded and texture-deficient regions, enhancing disparity estimation where traditional matching fails (see Figure 4 and 6).
Foundational Contribution to Vision Understanding: deepens our understanding of 3D scene geometry, paving the way for more interpretable and reliable vision systems.
@@ 2.3 Occlusions, Discontinuities and GCs
Following Proposition 2.4 we model surfaces $\mathcal{S}=\{(e,\mathcal{P}_{e});e=0,\ldots,M-1\}$ as a set of paths $\mathcal{P}_{e}$ in the cyclopean space and indexed by each epipolar line $e$. A path $\mathcal{P}_{e}=\{[O(e,x),d(e,x)];\,x=0,\frac{1}{2},1,\frac{3}{2},\ldots N-1,N-\frac{1}{2}\}$ is described by a sequence of states $[O(e,x),d(e,x)]$ for each $(e,x)$, where a binary occlusion variable $O(e,x)=0,1$ and the disparity value $d(e,x)$ are unique (that is, satisfying $GC2$). Here we do not distinguish whether a surface is described in terms of depth or disparity values. In order to impose $GC1$ the path $\mathcal{P}_{e}$ obeys the following constraint
@@
$1=O(e,x+\frac{1}{2})=O(e,x+1)=\ldots=O(e,x^{\prime}-\frac{1}{2})$
@@
$O(e,x)O(e,x-\frac{1}{2})=1\rightarrow d(e,x)-d(e,x-\frac{1}{2})=\pm\frac{1}{2}\,.$ (5)
Note that the disparity values assigned at such occlusion locations are not part of the final solution, they are just placeholders to insure that $GC1$ is satisfied. This specific characterization in XD, to the best of our knowledge, has not been proposed before, and it is a critical description to impose $GC1$ locally. Later, the disparity values at the occlusions will have to be estimated.
@@ 2.4 Optimization Criteria and DP
${\bf C}({\mathcal{P}_{e}})=\sum_{x}\lambda\,O(e,x)-\epsilon\,O(e,x)\,O(e,x-\frac{1}{2})\hskip 3.61371pt$
@@
where the hyperparameters $\lambda,\epsilon$ estimates, for a given path, if the quality of the match between features at each $(e,x)$ is better than assigning to it an occlusion $O(e,x)=1$. The lower $\lambda$ is the better the match must be in order not to be chosen as an occlusion. The higher $\epsilon$, the more likely $O(e,x)O(e,x\pm\frac{1}{2})=1$.
@@
=$=\left\{[O(e,x-\frac{1}{2})=0,1;d(e,x-\frac{1}{2})=d(e,x)\pm 0,\frac{1}{2}\right\}\,.$
@@ 2.5 Detection of Homogeneous Regions
The detection of homogeneous regions – regions with a lack of texture – occurs during the detection of occlusion in DP. The assignment of occlusions in homogeneous regions is caused by lack of a good match (they are equally bad matches) stereo features in homogeneous regions. More precisely, we simply assign $h(e,x)=1$ and $h(e,x-\frac{1}{2})=1$ to locations $(e,x),(e,x-\frac{1}{2})$ where DP assigned $O(e,x)=O(e,x-\frac{1}{2})=1$ and yet (5) is not satisfied, i.e., where also $0=d(e,x)-d(e,x-\frac{1}{2})$. An illustrative example of the results of DP is shown in Figure 5 Step (8). Note that the disparity values at the occlusions and homogeneous regions are not reliable values and will still have to be estimated.
For space reasons, we do not elaborate here further on the exact details of the DP algorithm; instead, we make it available on Github [^1] https://github.com/SherlonAlmeida/B2FS.git.
@@ 3 Back to the Future
The model we developed in section 2 focused on geometric reasoning, a set of abstractions about the 3D scene as viewed from the cyclopean eye. The DP algorithm employed is a method to obtain occlusions and homogeneous regions as well as disparity (depth) in visible regions (where the stereo features ${F^{L},F^{R}}$ do match well) with subpixel accuracy. We still must obtain such features ${F^{L},F^{R}}$, and invoke a surface model to fill in depth at occlusions and homogeneous regions, and possibly improve the accuracy of the depth obtained from DP to a decimal value. We refer to this approach as B2FS and it is depicted in Figure 5 with all steps of B2FS shown and described. We now describe the steps in more depth and at a high level, pointing to the figures and their captions for more detail information.
Figure 5: Workflow. Step (1) our approach receives an image-pair with resolution $r^{H,W,3}$. In order to be obtain full resolution features and in anticipation of RAFT-Stereo resolution reduction, Step (2) upscale to $r^{4H,4W,3}$ using Hybrid Attention Transformer [5]. Step (3), RAFT-Stereo performs at $\frac{1}{4}$ of $r^{4H,4W,3}$, resulting in $r^{H,W,256}$ for ${F^{L},F^{R}}$. Step (4) uses bilinear interpolation to produce twice as much the width resolution $r^{H,2W,256}$, and achieve the subpixel data for the XD space (the red dots in Figure 2). Steps (5) and (6) perform the feature dot products to create the correlation matrix by epipolar lines at resolution $r^{2W,2W,1}$. Step (7) transfer the correlation data to the XD space, with the max disparity considered shown by a solid red horizontal line (cutting the need to search beyond such disparity). Step (8) run DP to obtain a Disparity mask at good data matching coordinates i.e., where a binary Data mask indicates the non occluded and non homogeneous regions. Step (8) also performs monocular depth. Step(9) performs the fill in information of the surface where DP did not have a solution (homogeneous and occlusions) via a FCRN where the input is a normalized monocular depth and the output is the DP disparity map only where such solution is available by the data mask.
@@ 3.1 Image Features for Stereo Matching
Considering the promising results from the DL methods already developed that extract stereo features from stereo image pairs, we adopted RAFT-Stereo [16] features as input to the DP program (see Figure 6). However, RAFT-Stereo produces features at $\frac{1}{4}$ of the input image resolution. An interpolation must be considered to achieve full resolution features and is shown in Figure 5, steps (6) and (7), leading to the LR correlation and data filling of the XD space. Note that Figure 6 already shows the LR correlation at full resolution (after interpolation).
Figure 6: LR Correlation: The graph displays the distance $FM(e,x,d(e,x))=1-\frac{FMS(e,x,d(e,x))}{max(FMS(e,x,d(e,x))}$, where $FM$ is obtained from (6). Dark regions (low values of $FM$) represent good matches. The disparity inverse relation to depth guarantees that disparities are always equal or greater than zero, avoiding the search for negative disparities. One observes that typically one disparity stands out (a good match), unless we have occlusions or homogeneous regions. One can identify regions of occlusions where (i) no good match along the disparity axis occur and where (ii) the occlusion width is similar to the disparity changes from neighbor dark segment, as proposed by constraint $GC1$.
@@ 3.2 Surface Model
Our approach to model surfaces, needed to complete the surfaces where occlusions and homogeneous regions are masked by the DP solution, relies on results in monocular depth using DL. Excellent depth gradient information is obtained with Depth Anything V2 [27] and DepthPro [3]. As we experimented, the depth values offered are not scaled to precision (see Figure 8 for the analysis). Nevertheless, the accuracy of the depth-gradient information, which gives the surface normal unit vectors, is in par with the decimal disparity accuracy we aim to obtain. Thus, we treat monocular depth solutions as the best models today for providing surface normals.
After conducting a study on Depth Anything V2 and DepthPro we observed that the range of the 3D environment captured for DepthPro is wider and edges more accurate than Depth Anything V2. In addition, it provides an extra feature: an estimation of focal length for a monocular input image, which allows us to estimate the disparity map from depth using the Equation (2). Note that monocular depth solutions provide surface information from the view of the provided image and not from the cyclopean eye view. The differences in such depth values are captured by Figure 3 and Equation (3).
@@
3.3 Combining Stereo x Mono Disparities
@@ 4 Results and Discussion
We evaluated our approach using the Middlebury dataset, which consists of realistic indoor scenes, and in this section we discuss our quantitative and qualitative findings for $r^{256,256,3}$ resolution images.
Initially, we compared the disparities estimated by DepthPro with the GT using the average error (AvgError). We obtained it based on the focal length estimated by DepthPro, and baseline provided by GT. Figures  (8, 9) demonstrate that monocular estimation produces out of scale disparities for all images. Our technique effectively leverage monocular gradients, feature matching and stereo reasoning about surfaces (B2FS) to obtain accurate solutions across all Middlebury images with public GT (Additional and Training partitions), reducing AvgError by up to 48 times.
@@
Figure 9: Quantitative Results: B2FS vs. DepthPro (Normalized-GT). In this experiment, we normalized DepthPro disparities between the minimum and maximum disparities from GT. It is evident that simply transforming DepthPro into the correct disparity range is insufficient to achieve optimal results, indicating that obtaining accurate disparities is not solely dependent on a scale factor.
We obtained the publicly available code for state-of-the-art techniques such as RAFT-Stereo [16] and Selective-IGEV [24] to assess their robustness in extracting stereo cues from input images with lower resolutions than those used in benchmarks such as Middlebury, ETH3D, and KITTI. Our study focused on images $r^{512,512,3}$, $r^{256,256,3}$, and $r^{128,128,3}$. As shown in Figure 10, RAFT-Stereo and Selective-IGEV struggle as the input resolution decreases, while B2FS maintains the surface appearance.
@@
Our evaluation includes standard benchmark metrics: AvgError, BadError, and RMSError. While our numerical results (see Figure 12) were not superior to RAFT-Stereo and Selective-IGEV, the qualitative results remained consistent across different resolutions. Additionally, Figure 11 shows that our approach exhibits more global errors than local errors relative to RAFT-Stereo and Selective-IGEV, indicating a tendency to preserve local details.
@@
Given that our qualitative results preserved detailed monocular surface information while maintaining DP GCs (see Figure 1), we questioned the effectiveness of AvgError, BadError, and RMSError in assessing the robustness of monocular and stereo techniques. To address this, we also evaluated the inverse of the Structural Similarity Index Measure (SSIMError, where lower is better), as well as Mutual Information Similarity (MutualInfoSim) and Peak Signal-to-Noise Ratio (PSNRSim), where higher values indicate better performance. Our results (blue) outperformed RAFT-Stereo (green) in most cases and were competitive with Selective-IGEV.
These findings highlight the limitations of traditional error metrics in evaluating monocular and stereo depth estimation techniques. While AvgError, BadError, and RMSError provide numerical accuracy assessments, they do not fully capture the perceptual quality or structural consistency of disparity maps. Our results demonstrate that B2FS preserves finer details and maintains structural coherence, particularly in lower-resolution inputs where stereo-based methods struggle. This suggests that leveraging stereo constraints and monocular gradients can offer a robust alternative, even when input resolutions deviate from those used in training. B2FS effectively balances depth accuracy with high-quality visual reconstruction, making it promising for real-world applications where resolution constraints are common.

### v2

# BACK TO THE FUTURE CYCLOPEAN STEREO: A Human Perception Approach Combining Deep and Geometric Constraints

@@ 1 Introduction
What does it mean to “understand stereo vision” and why does it matter? Understanding means discovering internal representations of the problem at hand, and often it is needed to generalize to new situations or to allow us to apply such knowledge broadly. Suppose that a camera records video data of objects falling to the ground. It is plausible that an algorithm from machine learning (ML) could make predictions of such trajectories with an accuracy equal to or greater than any physics model, assuming enough variety and quantity of data are provided. However, such an ML solution would not have revealed the underlying laws of motion that govern them, such as the formula $F=ma$ and $a=9.8m/s^{2}$, an abstraction of reality that alone does not even give the most accurate models when the wind plays a role in the trajectory. Still, it has been a key insight for engineering and our civilization. Our goal is to better understand stereo vision with the mandate that to demonstrate better understanding, the new solution must perform on par or better with respect to the state-of-the-art stereo vision. In this way, the abstractions of the 3D reality we make here can help not only to improve efficiency and performance of stereo algorithms, but perhaps such 3D models of reality can be used for other activities in computer vision.
@@ 1.2 Our Contribution
Cyclopean Eye Model: a novel framework that extracts depth discontinuities and occlusions, imposing a human-like perception of 3D scenes that for opaque surfaces there is one and only one disparity solution per cyclopean spatial coordinate $(e,x)$ (see Figures 2, 3, and 4).
Occlusion and Texture-Less Region Depth Recovery: incorporate a prior monocular surface model to fill in the depth values where occluded and texture-deficient regions occur (see Figures 4 and 6).
Foundational Contribution to Vision Understanding: deepens our understanding of 3D scene geometry, paving the way for more interpretable, generalizable, and reliable vision systems.
@@ 2.3 Occlusions, Discontinuities and GCs
Following Proposition 2.4 we model surfaces $\mathcal{S}=\{(e,\mathcal{P}_{e});e=0,\ldots,M-1\}$ as a set of paths $\mathcal{P}_{e}$ in the cyclopean space and indexed by each epipolar line $e$. A path $\mathcal{P}_{e}=\{[O(e,x),d(e,x)];\,x=0,1/2,1,3/2,\ldots N-1,N-1/2\}$ is described by a sequence of states $[O(e,x),d(e,x)]$ for each $(e,x)$, where a binary occlusion variable $O(e,x)=0,1$ and the disparity value $d(e,x)$ are unique (that is, satisfying $GC2$). Here we do not distinguish whether a surface is described in terms of depth or disparity values. In order to impose $GC1$ the path $\mathcal{P}_{e}$ obeys the following constraint
@@
$1=O(e,x+1/2)=O(e,x+1)=\ldots=O(e,x^{\prime}-1/2)$
@@
$O(e,x)O\left(e,x-\frac{1}{2}\right)=1\rightarrow d(e,x)-d\left(e,x-\frac{1}{2}\right)=\pm\frac{1}{2}\,.$ (5)
Note that the disparity values assigned at such occlusion locations are not part of the final solution, they are just placeholders to ensure that $GC1$ is satisfied. This specific characterization in XD, to the best of our knowledge, has not been proposed before, and it is a critical description to impose $GC1$ locally. Later, the disparity values at the occlusions will have to be estimated.
@@ 2.4 Optimization Criteria and DP
${\bf C}({\mathcal{P}_{e}})=\sum_{x}\lambda\,O(e,x)-\epsilon\,O(e,x)\,O(e,x-1/2)\hskip 3.61371pt$
@@
where the hyperparameters $\lambda,\epsilon$ estimates, for a given path, if the quality of the match between features at each $(e,x)$ is better than assigning to it an occlusion $O(e,x)=1$. The lower $\lambda$ is the better the match must be in order not to be chosen as an occlusion. The higher $\epsilon$, the more likely $O(e,x)O(e,x\pm 1/2)=1$.
@@
=$=\left\{[O(e,x-1/2)=0,1;d(e,x-1/2)=d(e,x)\pm 0,1/2\right\}\,.$
@@ 2.5 Detection of Homogeneous Regions
The detection of homogeneous regions—areas lacking texture—occurs during occlusion detection in DP. Occlusions in homogeneous regions arise due to the absence of a good match, as all stereo features in these regions are equally poor matches. More precisely, we simply assign $h(e,x)=1$ and $h(e,x-\frac{1}{2})=1$ to locations $(e,x),(e,x-\frac{1}{2})$ where DP assigned $O(e,x)=O(e,x-\frac{1}{2})=1$ and yet (5) is not satisfied, i.e., where also $0=d(e,x)-d(e,x-\frac{1}{2})$. An illustrative example of the results of DP is shown in Figure 5 Step (8). Note that the disparity values at the occlusions and homogeneous regions are not reliable values and will still have to be estimated.
We make the DP available on Github [^1] https://github.com/SherlonAlmeida/B2FS.git.
@@ 3 Back to the Future
The model we developed in section 2 focused on geometric reasoning, a set of abstractions about the 3D scene as viewed from the cyclopean eye. The DP algorithm employed is a method to obtain occlusions and homogeneous regions as well as disparity (depth) in visible regions (where the stereo features ${F^{L},F^{R}}$ do match well) with subpixel accuracy. We still must obtain such features ${F^{L},F^{R}}$, and invoke a surface model to fill in depth at occlusions and homogeneous regions, and possibly improve the accuracy of the depth obtained from DP to a decimal value. We refer to this approach as B2FS and it is depicted in Figure 5 with all steps of B2FS shown and described. We now describe the steps in more depth and at a high level, pointing to the figures and their captions for more detailed information.
Figure 5: Workflow. Step (1) our approach receives an image-pair with resolution $r^{H,W,3}$. In order to obtain full resolution features and in anticipation of RAFT-Stereo resolution reduction, Step (2) upscale to $r^{4H,4W,3}$ using Hybrid Attention Transformer [5]. Step (3), RAFT-Stereo performs at $\frac{1}{4}$ of $r^{4H,4W,3}$, resulting in $r^{H,W,256}$ for ${F^{L},F^{R}}$. Step (4) uses bilinear interpolation to produce twice as much the width resolution $r^{H,2W,256}$, and achieve the subpixel data for the XD space (the red dots in Figure 2). Steps (5) and (6) perform the feature dot products to create the correlation matrix by epipolar lines at resolution $r^{2W,2W,1}$. Step (7) transfer the correlation data to the XD space, with the max disparity considered shown by a solid red horizontal line (cutting the need to search beyond such disparity). Step (8) run DP to obtain a Disparity mask at good data matching coordinates i.e., where a binary Data mask indicates the non occluded and non homogeneous regions. Step (8) also performs monocular depth. Step (9) performs the fill in information of the surface where DP did not have a solution (homogeneous and occlusions) via a FCRN where the input is a normalized monocular depth and the output is the DP disparity map only where such solution is available by the data mask.
@@ 3.1 Image Features for Stereo Matching
Considering the promising results from the DL methods already developed that extract stereo features from stereo image pairs, we adopted RAFT-Stereo [16] features as input to the DP program (see Figure 6). However, RAFT-Stereo produces features at $\frac{1}{4}$ of the input image resolution. An interpolation must be added to achieve full resolution features and is shown in Figure 5, steps (6) and (7), leading to the LR correlation and data filling of the XD space. Note that Figure 6 already shows the LR correlation at full resolution (after interpolation).
Figure 6: LR Correlation: The graph displays the distance $FM(e,x,d(e,x))=1-\frac{FMS(e,x,d(e,x))}{max(FMS(e,x,d(e,x))}$, where $FMS$ is obtained from (6). Dark regions (low values of $FM$) represent good matches. The disparity inverse relation to depth guarantees that disparities are always equal or greater than zero, avoiding the search for negative disparities. One observes that typically one disparity stands out (a good match), unless we have occlusions or homogeneous regions. One can identify regions of occlusions where (i) no good match along the disparity axis occur and where (ii) the occlusion width is similar to the disparity changes from neighbor dark segment, as proposed by constraint $GC1$.
@@ 3.2 Surface Model
Our approach to a prior model surfaces to complete the depth at occlusions and homogeneous regions, relies on results in monocular depth using DL. Excellent depth gradient information is obtained with Depth Anything V2 [27] and DepthPro [3]. As we experimented, the depth values offered are not scaled to precision (see Figure 8 for the analysis). Nevertheless, the accuracy of the depth-gradient information, which gives the surface normal unit vectors, is in par with the decimal disparity accuracy we aim to obtain.
After conducting a study on Depth Anything V2 and DepthPro we observed that the range of the 3D environment captured for DepthPro is wider and edges more accurate than Depth Anything V2. Note that monocular depth solutions provide surface information from the view of the provided image and not from the cyclopean eye view. The differences in such depth values are captured by Figure 3 and Equation (3).
@@
3.3 Combining Stereo x Monocular Depth
@@ 4 Results and Discussion
We evaluated our approach using the Middlebury dataset, which consists of realistic indoor scenes. Our analysis includes both quantitative and qualitative findings, primarily for images with $r^{256,256,3}$ resolution. This resolution was chosen to balance quality and practicality—high enough for meaningful results and real applications (e.g., virtual reality) but not so high that thorough experimentation becomes impractical for our limited academic computing resources. Additionally, since most algorithms are designed for high-resolution images, lower resolutions pose a greater challenge for purely data-driven techniques, highlighting the need for the B2FS approach.
Several comparisons are provided to illustrate B2FS’s performance.
4.1 B2FS vs Monocular Depth
Monocular depth methods ”claim to be accurate and they do look good”. In order to compare DepthPro depth estimates with the GT disparities, we scale the depth estimates based on the baseline provided by GT and the focal length (either provided by Depth Pro or by the GT). Figure 8 demonstrate that monocular estimation produces out-of-scale disparities for all images. In contrast, B2FS obtains accurate solutions across all Middlebury images with public GT (Additional and Training partitions), reducing AvgError by up to 48 times.
We then examined the hypothesis of using the best possible affine transformation (translation and scaling) to the monocular depth that would minimize the error with respect to the GT, i.e., we used the GT to estimate the best affine transformation to be applied to monocular depth. Figure 9 shows that still B2FS outperforms such result.
@@ 4.1 B2FS vs Monocular Depth
Figure 9: Quantitative Results: B2FS vs DepthPro (Normalized-GT). In this experiment, using GT data, we applied an affine transformation (translation and scaling) to the DepthPro disparities, guaranteeing it will produce the minimum and maximum disparities provided by GT. Still B2FS outperform DepthPro, indicating that DepthPro up to an affine transformation can not obtain accurate disparities.
4.2 B2FS vs RAFT-Stereo, Selective-IGEV
We obtained the publicly available code for state-of-the-art techniques RAFT-Stereo [16] and Selective-IGEV [24] to assess their robustness in extracting stereo cues from input images with lower resolutions than those used in benchmarks such as Middlebury, ETH3D, and KITTI.
One study focused on images with resolutions $r^{512,512,3}$, $r^{256,256,3}$, and $r^{128,128,3}$. Figure 10 shows that RAFT-Stereo and Selective-IGEV struggle as the input resolution decreases, while B2FS maintains the surface appearance. Here, we emphasize that B2FS is not overly trained on any particular resolution.
@@ 4.2 B2FS vs RAFT-Stereo, Selective-IGEV
Focusing in resolution $r^{256,256,3}$ and standard benchmark with metrics: AvgError, BadError, and RMSError, our numerical results (see Figure 12) are not superior to RAFT-Stereo and Selective-IGEV. In close inspection Figure 11 shows that our approach exhibits small errors everywhere but not large errors in high depth frequency locations relative to RAFT-Stereo and Selective-IGEV, indicating a tendency to our approach to preserve local depth structures.
@@
Given that our qualitative results preserved detailed surface information while maintaining DP GCs (see Figure 1), we questioned the effectiveness of AvgError, BadError, and RMSError in assessing the robustness of depth techniques. To address this, we also evaluated the inverse of the Structural Similarity Index Measure (SSIMError, where lower is better), as well as Mutual Information Similarity (MutualInfoSim) and Peak Signal-to-Noise Ratio (PSNRSim), where higher values indicate better performance (see Figure 12). Our results (blue) outperformed RAFT-Stereo (green) in most cases and were competitive with Selective-IGEV.
These findings highlight the limitations of traditional error metrics in evaluating depth estimation techniques. While AvgError, BadError, and RMSError provide numerical accuracy assessments, they do not fully capture the perceptual quality or structural consistency of disparity maps. Our results demonstrate that B2FS preserves finer details and maintains structural coherence, particularly in lower-resolution inputs where stereo-based methods struggle. This suggests that leveraging stereo geometric constraints and monocular gradients can offer a robust alternative, even when input resolutions deviate from those used in training. B2FS effectively balances depth accuracy with high-quality visual reconstruction, making it promising for real-world applications where resolution constraints are common.
@@ 5 Conclusion
Acknowledgements
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brasil (CAPES) – Finance Code 001.
Supplementary Material
6 Overview
In this document, we extend the demonstration of the main contributions of the B2FS approach, emphasizing a visual presentation.
In Section 7 we revisit the two proposed geometric constraints $GC1$ and $GC2$ from direct data visualization as well as an inspection of occlusions, homogeneous regions, and data matching.
In Section 8 we provide visualization for the qualitative comparison of maintaining details between RAFT-Stereo, Selective-IGEV, and B2FS for all the 13 images of the additional partition of Middlebury dataset: Backpack, Bicycle, Cable, Classroom1, Couch, Flowers, Mask, Shopvac, Sticks, Storage, Sword1, Sword2, and Umbrella.
In Section 9 a visual comparison of the disparity errors per pixel suggests that B2FS performs significantly better at depth discontinuities, but slightly underperformed at homogeneous regions due to B2FS discarding feature matching there. Since homogeneous regions are large regions of the image, B2FS end up with more total error. Future work on integrating feature matching in these regions could improve the results. All images presented in Section 8 and Section 9 have a resolution of $r^{256,256,3}$ and include a caption that specifies the image of the data set used.
Finally, we present in Section 10 a study about the impact on the results of changing image resolution, and Section 11 concludes this document with general comments.
7 Matching, Occlusions and Homogeneity
This section examines data to validate the geometric reasoning presented by Proposition 2.4, constraints $GC1$ and $GC2$, and to validate the challenge of feature matching in homogeneous regions. See Figures 1314, and 15 and their respective captions.
Figure 13: The LR space displays $FM$ distances associated with the epipolar lines $e=128$ (yellow), where $FM_{e,x}(d)\in[0,1]$. Dark regions represent low $FM$ and good matches. $GC2$ is verified with R-occlusions in dashed-red, no matches, associated with the L-discontinuity (vertical jumps) and L-occlusions in dashed-blued, no matches, associated with R-discontinuity (horizontal jumps). From the cyclopean eye these are 45 degrees changes and $GC1$, unique disparity, is also verified.
Figure 14: The LR space displays $FM$ distances associated with the epipolar lines $e=128$ (yellow), where $FM_{e,x}(d)\in[0,1]$. Dark regions represent low $FM$ and good matches. $GC2$ is verified with R-occlusions in dashed-red, no matches, associated with the L-discontinuity (vertical jumps) and L-occlusions in dashed-blued, no matches, associated with R-discontinuity (horizontal jumps). From the cyclopean eye these are 45 degrees changes. In particular in order for stereo to output the front post depth (disparity) $GC1$, a unique disparity constraint, will have to be imposed to discard other possible good feature matches (low $FM$ values) as shown in the right figure.
Figure 15: Homogeneity: The LR space displays $FM$ distances associated with the epipolar lines $e=30$ (yellow), where $FM_{e,x}(d)\in[0,1]$. Dark regions represent low $FM$ and good matches. Note that shadows do yield good matches. However, in order to fill in homogeneous regions, surface priors become necessary. In our approach, we detect such homogeneous regions and discard feature matching values at these regions.
8 Qualitative results: a visual analysis
For each image, we selected a green rectangle to highlight a region of interest and zoomed in for better visualization. The results support the robustness of B2FS in preserving the scene structure. See figures 16171819202122232425262728.
Figure 16: Backpack
Figure 17: Bicycle
Figure 18: Cable
Figure 19: Classroom1
Figure 20: Couch
Figure 21: Flowers
Figure 22: Mask
Figure 23: Shopvac
Figure 24: Sticks
Figure 25: Storage
Figure 26: Sword1
Figure 27: Sword2
Figure 28: Umbrella
9 Quantitative visual results: signed error
We show the signed error for each pixel across the entire image, using red to indicate regions where the disparities should be further back (negative difference values) and blue for regions where disparities should be further ahead (positive difference values). White indicates areas where the error is not measured due to limitations in Middlebury disparity estimation. The signed error value at the bottom of each image is the average value of the absolute pixel signed error. Although B2FS has larger overall signed error, it is typically due to a slightly larger error estimation of depth in large homogeneous regions, where feature matching was discarded. At depth discontinuities, B2FS performs significantly better. Future work includes efforts to integrate homogeneous regions feature matching to guide monocular depth solutions. See figures 29303132333435363738394041.
Figure 29: Backpack
Figure 30: Bicycle
Figure 31: Cable
Figure 32: Classroom1
Figure 33: Couch
Figure 34: Flowers
Figure 35: Mask
Figure 36: Shopvac
Figure 37: Sticks
Figure 38: Storage
Figure 39: Sword1
Figure 40: Sword2
Figure 41: Umbrella
10 Impact of image resolution
In this section, the figures present the impact of resolution changes in the input images for RAFT-Stereo, Selective-IGEV, and B2FS. Notice that as the input image resolution decreases, B2FS preserves details, while RAFT-Stereo and Selective-IGEV struggle. We hypothesize that this effect is due to the DL algorithms being overly focused on higher-resolution images.
11 General Comments
Data-driven models rely on the quality and representativeness of data to accurately capture the properties of the target domain. Currently, the most accurate stereo vision techniques are trained on thousands of synthetic and/or real image pairs to generate high-quality disparity maps. Many studies in the literature focus on creating increasingly large datasets to train deep learning (DL) models, enhancing their performance. With sufficient data coverage, zero-shot scenarios become less common as models are exposed to a vast range of situations.
While expanding datasets is essential for faithfully capturing real-world patterns, it is not the only step necessary for understanding them. It is important to emphasize that purely data-driven models can also incorporate a reasoning step based on world priors related to the task. This can lead to more general results and simpler models, going beyond data fitting. See figures 42434445464748495051525354.
Figure 42: Backpack
Figure 43: Bicycle
Figure 44: Cable
Figure 45: Classroom1
Figure 46: Couch
Figure 47: Flowers
Figure 48: Mask
Figure 49: Shopvac
Figure 50: Sticks
Figure 51: Storage
Figure 52: Sword1
Figure 53: Sword2
Figure 54: Umbrella